OpenStudy (goformit100):

Please Help! Given A={-1,0,2,5,6,11}, B={-2,-1,0,18,108} and f(x)=x^2-x-2. Is f(A)=B? Find f(A).

OpenStudy (campbell_st):

Substitute each value in A to see if its in B f(-1) = (-1)^2 -(-1) -2 f(-1) = 0 f(0) = (0)^2 - (0) - 2 f(0) = -2 f(2) = (2)^2 - (2) - 2 = 0 etc just repeat it for each value in A

OpenStudy (anonymous):

plug in each value from the set of A, so f(A) and if the result is in the set of corresponding B, then f(A) = B ^^ like campbell said

OpenStudy (goformit100):

thankyou @javawarrior and @campbell_st